Use Rolle's theorems to show that there is a solution of the equation cotx=x in ]0,π/2[


1
Expert's answer
2021-01-05T14:58:08-0500

Let:

"f(x)=cosx-xsinx"

This function is continuous on the interval "[0,\pi/2]"

"f(0)=1, f(\pi/2)=-\pi/2"

So, by Intermediate Value Theorem there is "c\isin[0,\pi/2]" , such that "f(c)=0"

That is, "c" is the solution of the given equation.